# Ming-Flash-Omni: A Sparse, Unified Architecture for Multimodal Perception and Generation

### v1

@@ 1 Introduction
Driven by advances in Large Language Models (LLMs) and extensive training on large-scale multimodal datasets, Multi-modal Large Language Models (MLLMs) have demonstrated remarkable perceptual capabilities in both vision (Chen et al., 2024d; Bai et al., 2025b; KimiTeam et al., 2025; Xu et al., 2025c) and audio (Ding et al., 2025a; Xu et al., 2025a, c), as well as generative capabilities in these two modalities (Huang et al., 2025; Ding et al., 2025a; OpenAI, 2025; Tong et al., 2024b; Pan et al., 2025; Xu et al., 2025c). Nevertheless, effectively integrating comprehension and generation across multiple modalities into a unified model remains challenging. While humans naturally learn by combining multiple modalities, leveraging their complementary strengths and interactions to enhance overall learning efficiency, building a unified Omni-MLLM is hindered by representational disparities and modality imbalances.
In this paper, we introduce Ming-Flash-Omni, which builds upon the Ming-Omni architecture with a redesigned foundation and targeted enhancements across multimodal understanding and generation. At its core, Ming-Flash-Omni adopts Ling-Flash-2.0 lin (2025) (a scaled-up, highly sparse Mixture-of-Experts architecture) where an increased sparsity ratio enables substantial model capacity while maintaining bounded inference latency, striking a favorable trade-off between performance and efficiency.
On the understanding side, the model introduces two key advances. First, Ming-Flash-Omni upgrade the positional encoding to VideoRoPE Wei et al. (2025), a refined variant specifically designed to better capture temporal dynamics in video sequences, thereby enhancing the model’s ability to understand complex visual events. Second, Ming-Flash-Omni focus on improving the context-aware ASR capability itself, enhancing the model’s ability to leverage surrounding linguistic context during speech recognition and thereby achieving more accurate transcription in context-dependent scenarios.
@@
These architectural innovations empower Ming-Flash-Omni to deliver exceptional cross-modal performance in both comprehension and generation tasks. Specifically, in the image perception task, Ming-Flash-Omni attained performance comparable to that of Qwen3-Omni (Xu et al., 2025c). Ming-Flash-Omni also delivers superior performance in end-to-end speech understanding and generation.For instance, it achieves SOTA on all 12 metrics on ContextASR-Bench.
@@ 2 Ming-Flash-Omni
As illustrated in Figure 2, Ming-Flash-Omni retains the unified two-stage pipeline of Ming-Omni AI et al. (2025a), where perception supports multimodal understanding and generation targets speech and image synthesis, while markedly advancing long-context modeling, reasoning, and controllable generation. At the core is Ling‑Flash‑2.0 lin (2025), a sparse MoE LM (100B; 6.1B per token) with a dual balancing scheme that stabilizes training and improves efficiency. On the perception side, Ming-Flash-Omni employs VideoRoPE Wei et al. (2025) for temporal modeling, context‑aware ASR for more reliable speech understanding, and a think mode for deeper multi‑step reasoning. On the generation side, we replace discrete speech tokens with continuous acoustic latents, avoiding quantization loss and improving fidelity; for images, we upgrade to a synergistic training paradigm that enables generative segmentation-as-editing, facilitating fine-grained and controllable generation. Overall, Ming-Flash-Omni advances the unified model with stable expert routing and scalable long‑context modeling, yielding more reliable multimodal understanding and controllable generation.
@@ 2.1 Unified Understanding Across Modalities
The cornerstone of Ming-Flash-Omni is enhanced multimodal understanding. We retain the established visual and audio encoders (Qwen2.5 (Bai et al., 2025a) and Whisper (Radford et al., 2023)) and feed their projected embeddings, concatenated with tokenized text, into Ling‑flash‑2.0, a sparse MoE language model with distinct routers per modality. Beyond this, Ming-Flash-Omni incorporates VideoRoPE to maintain temporal coherence over long-range frame sequences, thus emphasizing temporal modeling. Furthermore, Ming-Flash-Omni adopt a context‑aware ASR training paradigm that conditions decoding on task or domain context, addressing common shortcomings of conventional ASR in real‑world, multi‑domain scenarios (limited world knowledge and unreliable proper‑noun recognition) and yielding more accurate, context‑consistent transcripts. To stabilize training in the more sparse Ling‑flash‑2.0, we employ a hybrid expert‑balancing scheme that combines an auxiliary load‑balancing loss (as in Ming-Omni) with per‑router bias updates (as in Ling‑flash‑2.0), promoting uniform expert activation and improved convergence.
@@ 2.2 Unified Speech Understanding and Generation
Ming-Flash-Omni retains the same audio encoder as Ming-Omni (AI et al., 2025a), with optimization efforts focused on improving contextual automatic speech recognition (ASR) performance by incorporating preceding textual context and hotword lists during training. For generation, we replace discrete speech tokens with continuous acoustic latents, avoiding quantization loss and improving fidelity. Specifically, we utilize a fixed, pre-trained audio head based on Qwen2.5 (0.5B parameters), which takes LLM-generated text tokens and downsampled VAE latents as input and autoregressively predicts the conditioning signals for the flow-matching head—following the paradigm of Jia et al. (2025).
It is worth noting that the continuous features used for our generation tasks are derived from our unified continuous speech tokenizer, which is based on a VAE-GAN architecture, with the overall training objective comprising two main components: a generator loss and a discriminator loss. The generator loss consists of (i) a multi-scale mel-spectrogram reconstruction loss, (ii) an adversarial loss, (iii) a feature matching loss, and (iv) a KL divergence regularization term. The discriminator, in turn, is trained primarily with the adversarial loss. For more implementation details, please refer to Canxiang et al. (2025).
@@ 2.4 Overall Training Procedure
The perception stage, consistent with the Ming-Omni, includes pre-training, instruction tuning, alignment tuning, and an additional coherent RL phase. The coherent RL phase proceeds sequentially: Dynamic‑GRPO (D-GRPO) AI et al. (2025b) followed by Unified-DPO (U-DPO). First, we use D‑GRPO, applying RL with verifiable rewards on datasets with checkable answers (via on‑policy GRPO with balanced sampling, dynamic hyperparameter adjustment, and task‑specific accuracy rewards) to reinforce correct reasoning across multimodal tasks. Next, we employ U‑DPO for preference alignment, extending standard DPO with an auxiliary instruction‑tuning loss term on the chosen samples to enhance instruction adherence and stylistic coherence. Concretely, we employ a replay strategy with instruction‑stage multimodal data and interleave updates on DPO and instruction‑tuning data to stabilize optimization and maintain balanced gradient flow.
After perception, we freeze the perception MLLM and optimize only the image generator, while leveraging the pre-trained audio generator from Canxiang et al. (2025). For image generation, the training procedure contains two sequential stages. In the first stage, we pre-train a diffusion-based image generator using a flow matching objective, while keeping the perception MLLM frozen. The generator is equipped with multi-scale learnable queries to capture hierarchical visual semantics from textual inputs. In the second stage, we extend the model to support image editing by conditioning the denoising process on reference images: the VAE-encoded representations of input images are concatenated with the noisy latent to enforce structural and semantic consistency with the original content. Additionally, input word-level captions are encoded via ByteT5 embeddings to enrich textual conditioning.
@@ 2.5 Infrastructure
To address these challenges, Ming-Flash-Omni is trained on an enhanced version of the Megatron-LM Shoeybi et al. (2019) framework with two key extensions tailored for multimodal workloads:
@@ 3 Data Construction
We have collected a large and diverse set of training data to enable models to process and understand information from multiple modalities, including text, images, audio and videos. The majority of this data comes from Ming-Omni (AI et al., 2025a). In addition, we develop several data processing pipelines to ensure data quality, diversity and deduplication. Establishing an effective multimodal data strategy is essential for the joint multi-modal training, as it facilitates seamless alignment of knowledge across diverse modalities. We categorize the training data based on the core modalities they are designed to enhance, including image, audio, video, and text. The detailed sources and construction methods for each type of data are elaborated in this section.
@@ 3.1 Image Data
Image data serves as the cornerstone of our multi-modal corpus. Following Ming-Omni (AI et al., 2025a), we integrate both image-understanding and image-generation datasets to enable the MLLM to acquire unified perception and generation capabilities. Additionally, we further design novel pipelines to synthesize high-quality datasets across diverse dimensions to improve model’s capabilities and user interaction quality.
@@ 3.1.1 Image Understanding Data
Reasoning Data: In the reasoning training of Ming-Flash-Omni, we enrich the CoT data to enhance the model’s reasoning capabilities. These data are primarily constructed around three key themes: mathematical logic, spatial reasoning, and GUI reasoning. We design an efficient CoT generation and filtering pipeline to construct high-quality, well-structured reasoning data that enhances the model’s multi-step reasoning capability. (1) We sample long CoT data using state-of-the-art multimodal reasoning models (e.g., Gemini (Gemini et al., 2023)) to build an initial CoT pool. (2) We evaluate the accuracy and quality of the synthesized CoT data and filter out the low-quality data. Based on this pipeline, we construct 1.5M multimodal long CoT samples, with a maximum length of 16K tokens. Besides, to overcome the limitations of the “direct answer” pattern in text-based QA data, we use reinforcement learning models to generate multi-step reasoning traces and final answers. Experimental results demonstrate that this data significantly improves Ming-Flash-Omni’s performance on complex reasoning tasks.
Preference alignment Data: To further enhance user interaction and response quality, we introduce preference alignment data, focusing on three main aspects: (1) Instruction intent understanding: To enhance the model’s ability to understand the true intent behind user instructions, we design a novel instruction intent reasoning paradigm where the model first determines if the instruction is clear. For clear instructions, it provides structured reasoning and final answers. For ambiguous ones, it infers the likely intent, generates clarification prompts, and aligns responses accordingly. By incorporating experience-alignment training, we significantly improve the model’s intent understanding and enhance the overall user experience. (2) Multi-turn conversation: Users often engage in repeated questioning on the same context. To maintain semantic consistency across turns, we decompose complex instructions into multi-turn conversations and generate high-quality responses using expert-annotated LLM pipelines (Laban et al., 2025). This ensures context retention and reduces performance degradation in multi-turn scenarios. (3) Complex multimodal instruction following: Users often provide sequential, interdependent commands. We design a multimodal instruction generation pipeline to generate SFT and DPO-style data to cover basic and complex instruction types (Zhang et al., 2024; Qian et al., 2024; Ding et al., 2025b).
@@ 3.1.2 Image Generation Data
Image generation data enhances MLLM capabilities beyond traditional image understanding tasks. In addition to the image generation data used in Ming-Omni (AI et al., 2025a), we specifically integrate image segmentation data, text generation data and portrait preservation data to further improve the user experience.
Image segmentation data: To improve the model’s generative segmentation capability, we construct two types of data: (1) We use the open-source referring segmentation datasets RefCOCO/+/g (Kazemzadeh et al., 2014; Mao et al., 2016) to construct image editing data. The original image serves as the reference, and binary masks highlight target regions with specified colors to create edited images. (2) For semantic and panoptic segmentation, samples are built from COCO-Panoptic (Lin et al., 2014; Kirillov et al., 2019), where each class or instance is assigned a unique color via a predefined colormap to generate edited images.
Portrait preservation data: The portrait preservation data consist of two data sources: (1) ID Photo Dataset: We collected and constructed 200k paired lifestyle-ID photos. And filter the data using four criteria, e.g., face similarity, face size and confidence, face angle and manual review. (2) Landmark Check-In Portrait Dataset: We collect 20K high-quality portraits from 225 landmarks.The original images serve as edited images, while using advanced segmentation model like SAM2 (Ravi et al., 2024) to segments the foreground person and places them onto 1,000 manually collected daily background scenes as pre-edited images. And then use LLM generate diverse prompts for each landmark.
Text generation data: We build a Chinese-English text generation dataset across three difficulty levels: (1) Monotonic background text rendering: Text is rendered directly by setting background color, font type, size, color, and position. (2) Text rendering on existing images: texts are rendered on suitable smooth regions obtained by Felzenszwalb algorithm. (3) Text-image integrated rendering: Using the SOTA LLMs to generate text rendering prompts, the advanced generation models (e.g., Qwen-image (Wu et al., 2025) and Nano Banana) are used for image generation, followed by OCR for consistency checks, resulting in a high-quality dataset.
@@ 3.2 Audio Data
For audio data, we mainly use the data from Ming-Omni (AI et al., 2025a). In addition, we incorporate the following three datasets to further enhance the model’s audio understanding and generation capabilities.
@@
TTS Data: The diversity of TTS data is essential for fully leveraging the pretrained language model’s capabilities in audio generation. In addition to the open-source data used in Ming-Omni, we develop a data generation pipeline to create large-scale TTS data. Specifically, (1) we crawl extensive audio data using keywords expanded from handcrafted seeds through domain-specific lexical variations, (2) apply VAD (Gao et al., 2023) to segment well-conditioned short clips, and (3) iteratively train an audio labeler—initially on high-quality data, then using its predictions to label the corpus and refine its accuracy. Based on the short audio clips and the audio labeler, we acquire a large number of high-quality audio clips with labels from different domains.

### v2

@@ 1 Introduction
Driven by advances in Large Language Models (LLMs) and extensive training on large-scale multimodal datasets, Multi-modal Large Language Models (MLLMs) have demonstrated remarkable perceptual capabilities in both vision (chen2024internvl; bai2025qwen2; team2025kimi; xu2025qwen3) and audio (ding2025kimi; xu2025qwen2; xu2025qwen3), as well as generative capabilities in these two modalities (huang2025step; ding2025kimi; chatgpt4o; tong2024metamorph; pan2025transfer; xu2025qwen3). Nevertheless, effectively integrating comprehension and generation across multiple modalities into a unified model remains challenging. While humans naturally learn by combining multiple modalities, leveraging their complementary strengths and interactions to enhance overall learning efficiency, building a unified Omni-MLLM is hindered by representational disparities and modality imbalances.
In this paper, we introduce Ming-Flash-Omni, which builds upon the Ming-Omni architecture with a redesigned foundation and targeted enhancements across multimodal understanding and generation. At its core, Ming-Flash-Omni adopts Ling-Flash-2.0 lingv2_flash (a scaled-up, highly sparse Mixture-of-Experts architecture) where an increased sparsity ratio enables substantial model capacity while maintaining bounded inference latency, striking a favorable trade-off between performance and efficiency.
On the understanding side, the model introduces two key advances. First, Ming-Flash-Omni upgrade the positional encoding to VideoRoPE wei2025videorope, a refined variant specifically designed to better capture temporal dynamics in video sequences, thereby enhancing the model’s ability to understand complex visual events. Second, Ming-Flash-Omni focus on improving the context-aware ASR capability itself, enhancing the model’s ability to leverage surrounding linguistic context during speech recognition and thereby achieving more accurate transcription in context-dependent scenarios.
@@
These architectural innovations empower Ming-Flash-Omni to deliver exceptional cross-modal performance in both comprehension and generation tasks. Specifically, in the image perception task, Ming-Flash-Omni attained performance comparable to that of Qwen3-Omni (xu2025qwen3). Ming-Flash-Omni also delivers superior performance in end-to-end speech understanding and generation.For instance, it achieves SOTA on all 12 metrics on ContextASR-Bench.
@@ 2 Ming-Flash-Omni
As illustrated in Figure 2, Ming-Flash-Omni retains the unified two-stage pipeline of Ming-Omni ai2025ming, where perception supports multimodal understanding and generation targets speech and image synthesis, while markedly advancing long-context modeling, reasoning, and controllable generation. At the core is Ling‑Flash‑2.0 lingv2_flash, a sparse MoE LM (100B; 6.1B per token) with a dual balancing scheme that stabilizes training and improves efficiency. On the perception side, Ming-Flash-Omni employs VideoRoPE wei2025videorope for temporal modeling, context‑aware ASR for more reliable speech understanding, and a think mode for deeper multi‑step reasoning. On the generation side, we replace discrete speech tokens with continuous acoustic latents, avoiding quantization loss and improving fidelity; for images, we upgrade to a synergistic training paradigm that enables generative segmentation-as-editing, facilitating fine-grained and controllable generation. Overall, Ming-Flash-Omni advances the unified model with stable expert routing and scalable long‑context modeling, yielding more reliable multimodal understanding and controllable generation.
@@ 2.1 Unified Understanding Across Modalities
The cornerstone of Ming-Flash-Omni is enhanced multimodal understanding. We retain the established visual and audio encoders (Qwen2.5 (bai2025qwen25vltechnicalreport) and Whisper (radford2023robust)) and feed their projected embeddings, concatenated with tokenized text, into Ling‑flash‑2.0, a sparse MoE language model with distinct routers per modality. Beyond this, Ming-Flash-Omni incorporates VideoRoPE to maintain temporal coherence over long-range frame sequences, thus emphasizing temporal modeling. Furthermore, Ming-Flash-Omni adopt a context‑aware ASR training paradigm that conditions decoding on task or domain context, addressing common shortcomings of conventional ASR in real‑world, multi‑domain scenarios (limited world knowledge and unreliable proper‑noun recognition) and yielding more accurate, context‑consistent transcripts. To stabilize training in the more sparse Ling‑flash‑2.0, we employ a hybrid expert‑balancing scheme that combines an auxiliary load‑balancing loss (as in Ming-Omni) with per‑router bias updates (as in Ling‑flash‑2.0), promoting uniform expert activation and improved convergence.
@@ 2.2 Unified Speech Understanding and Generation
Ming-Flash-Omni retains the same audio encoder as Ming-Omni (ai2025ming), with optimization efforts focused on improving contextual automatic speech recognition (ASR) performance by incorporating preceding textual context and hotword lists during training. For generation, we replace discrete speech tokens with continuous acoustic latents, avoiding quantization loss and improving fidelity. Specifically, we utilize a fixed, pre-trained audio head based on Qwen2.5 (0.5B parameters), which takes LLM-generated text tokens and downsampled VAE latents as input and autoregressively predicts the conditioning signals for the flow-matching head—following the paradigm of jia2025ditar.
It is worth noting that the continuous features used for our generation tasks are derived from our unified continuous speech tokenizer, which is based on a VAE-GAN architecture, with the overall training objective comprising two main components: a generator loss and a discriminator loss. The generator loss consists of (i) a multi-scale mel-spectrogram reconstruction loss, (ii) an adversarial loss, (iii) a feature matching loss, and (iv) a KL divergence regularization term. The discriminator, in turn, is trained primarily with the adversarial loss. For more implementation details, please refer to MingUniAudio2025.
@@ 2.4 Overall Training Procedure
The perception stage, consistent with the Ming-Omni, includes pre-training, instruction tuning, alignment tuning, and an additional coherent RL phase. The coherent RL phase proceeds sequentially: Dynamic‑GRPO (D-GRPO) ai2025m2 followed by Unified-DPO (U-DPO). First, we use D‑GRPO, applying RL with verifiable rewards on datasets with checkable answers (via on‑policy GRPO with balanced sampling, dynamic hyperparameter adjustment, and task‑specific accuracy rewards) to reinforce correct reasoning across multimodal tasks. Next, we employ U‑DPO for preference alignment, extending standard DPO with an auxiliary instruction‑tuning loss term on the chosen samples to enhance instruction adherence and stylistic coherence. Concretely, we employ a replay strategy with instruction‑stage multimodal data and interleave updates on DPO and instruction‑tuning data to stabilize optimization and maintain balanced gradient flow.
After perception, we freeze the perception MLLM and optimize only the image generator, while leveraging the pre-trained audio generator from MingUniAudio2025. For image generation, the training procedure contains two sequential stages. In the first stage, we pre-train a diffusion-based image generator using a flow matching objective, while keeping the perception MLLM frozen. The generator is equipped with multi-scale learnable queries to capture hierarchical visual semantics from textual inputs. In the second stage, we extend the model to support image editing by conditioning the denoising process on reference images: the VAE-encoded representations of input images are concatenated with the noisy latent to enforce structural and semantic consistency with the original content. Additionally, input word-level captions are encoded via ByteT5 embeddings to enrich textual conditioning.
@@ 2.5 Infrastructure
To address these challenges, Ming-Flash-Omni is trained on an enhanced version of the Megatron-LM shoeybi2019megatron framework with two key extensions tailored for multimodal workloads:
@@ 3 Data Construction
We have collected a large and diverse set of training data to enable models to process and understand information from multiple modalities, including text, images, audio and videos. The majority of this data comes from Ming-Omni (ai2025ming). In addition, we develop several data processing pipelines to ensure data quality, diversity and deduplication. Establishing an effective multimodal data strategy is essential for the joint multi-modal training, as it facilitates seamless alignment of knowledge across diverse modalities. We categorize the training data based on the core modalities they are designed to enhance, including image, audio, video, and text. The detailed sources and construction methods for each type of data are elaborated in this section.
@@ 3.1 Image Data
Image data serves as the cornerstone of our multi-modal corpus. Following Ming-Omni (ai2025ming), we integrate both image-understanding and image-generation datasets to enable the MLLM to acquire unified perception and generation capabilities. Additionally, we further design novel pipelines to synthesize high-quality datasets across diverse dimensions to improve model’s capabilities and user interaction quality.
@@ 3.1.1 Image Understanding Data
Reasoning Data: In the reasoning training of Ming-Flash-Omni, we enrich the CoT data to enhance the model’s reasoning capabilities. These data are primarily constructed around three key themes: mathematical logic, spatial reasoning, and GUI reasoning. We design an efficient CoT generation and filtering pipeline to construct high-quality, well-structured reasoning data that enhances the model’s multi-step reasoning capability. (1) We sample long CoT data using state-of-the-art multimodal reasoning models (e.g., Gemini (GEMINI)) to build an initial CoT pool. (2) We evaluate the accuracy and quality of the synthesized CoT data and filter out the low-quality data. Based on this pipeline, we construct 1.5M multimodal long CoT samples, with a maximum length of 16K tokens. Besides, to overcome the limitations of the “direct answer” pattern in text-based QA data, we use reinforcement learning models to generate multi-step reasoning traces and final answers. Experimental results demonstrate that this data significantly improves Ming-Flash-Omni’s performance on complex reasoning tasks.
Preference alignment Data: To further enhance user interaction and response quality, we introduce preference alignment data, focusing on three main aspects: (1) Instruction intent understanding: To enhance the model’s ability to understand the true intent behind user instructions, we design a novel instruction intent reasoning paradigm where the model first determines if the instruction is clear. For clear instructions, it provides structured reasoning and final answers. For ambiguous ones, it infers the likely intent, generates clarification prompts, and aligns responses accordingly. By incorporating experience-alignment training, we significantly improve the model’s intent understanding and enhance the overall user experience. (2) Multi-turn conversation: Users often engage in repeated questioning on the same context. To maintain semantic consistency across turns, we decompose complex instructions into multi-turn conversations and generate high-quality responses using expert-annotated LLM pipelines (laban2025llms). This ensures context retention and reduces performance degradation in multi-turn scenarios. (3) Complex multimodal instruction following: Users often provide sequential, interdependent commands. We design a multimodal instruction generation pipeline to generate SFT and DPO-style data to cover basic and complex instruction types (zhang2024iopo; qian2024mia; ding2025mm).
@@ 3.1.2 Image Generation Data
Image generation data enhances MLLM capabilities beyond traditional image understanding tasks. In addition to the image generation data used in Ming-Omni (ai2025ming), we specifically integrate image segmentation data, text generation data and portrait preservation data to further improve the user experience.
Image segmentation data: To improve the model’s generative segmentation capability, we construct two types of data: (1) We use the open-source referring segmentation datasets RefCOCO/+/g (refcoco; refcocog) to construct image editing data. The original image serves as the reference, and binary masks highlight target regions with specified colors to create edited images. (2) For semantic and panoptic segmentation, samples are built from COCO-Panoptic (Eval_mscoco; kirillov2019panoptic), where each class or instance is assigned a unique color via a predefined colormap to generate edited images.
Portrait preservation data: The portrait preservation data consist of two data sources: (1) ID Photo Dataset: We collected and constructed 200k paired lifestyle-ID photos. And filter the data using four criteria, e.g., face similarity, face size and confidence, face angle and manual review. (2) Landmark Check-In Portrait Dataset: We collect 20K high-quality portraits from 225 landmarks.The original images serve as edited images, while using advanced segmentation model like SAM2 (ravi2024sam) to segments the foreground person and places them onto 1,000 manually collected daily background scenes as pre-edited images. And then use LLM generate diverse prompts for each landmark.
Text generation data: We build a Chinese-English text generation dataset across three difficulty levels: (1) Monotonic background text rendering: Text is rendered directly by setting background color, font type, size, color, and position. (2) Text rendering on existing images: texts are rendered on suitable smooth regions obtained by Felzenszwalb algorithm. (3) Text-image integrated rendering: Using the SOTA LLMs to generate text rendering prompts, the advanced generation models (e.g., Qwen-image (wu2025qwen) and Nano Banana) are used for image generation, followed by OCR for consistency checks, resulting in a high-quality dataset.
@@ 3.2 Audio Data
For audio data, we mainly use the data from Ming-Omni (ai2025ming). In addition, we incorporate the following three datasets to further enhance the model’s audio understanding and generation capabilities.
@@
TTS Data: The diversity of TTS data is essential for fully leveraging the pretrained language model’s capabilities in audio generation. In addition to the open-source data used in Ming-Omni, we develop a data generation pipeline to create large-scale TTS data. Specifically, (1) we crawl extensive audio data using keywords expanded from handcrafted seeds through domain-specific lexical variations, (2) apply VAD (gao2023funasrfundamentalendtoendspeech) to segment well-conditioned short clips, and (3) iteratively train an audio labeler—initially on high-quality data, then using its predictions to label the corpus and refine its accuracy. Based on the short audio clips and the audio labeler, we acquire a large number of high-quality audio clips with labels from different domains.